The perimeter of the rectangle below is 48 units. The length of the rectangle is three times the width. What is the measurement of the length of this rectangle? The length = units.
Answer by addingup(3677)

2 x Length + 2 x Width = Perimeter. We can rewrite it like this:
2L + 2W = 48
The problem says that:
L = 3W
Let's plug this value for L in the formula above:
2(3W) + 2W = 48
6W + 2W = 48
8W = 48
W = 48/8 = 6
So, the width is 6 and the length is 3W = 3(6) = 18